After Larry had taken Nexium for 4 weeks, the pH in his stomach was raised to 4.52. What is the [H3O+] in his stomach?

Express the concentration to two significant figures and include the appropriate units.
Answer #1

pH = 4.52 pH – lo chot] -PH. [Hot] = 10 M - 4.52 M so (140+] = 10 -5 = 3.0199 X 10 M [Hot]= 3.0x105m